# Quantum Localization in Incommensurate Tight-Binding Chains

### v1

@@ Acknowledgements.
Appendix A “Almost decoupled” chains
@@
Let us set the inter-chain separation $d$ to a sufficiently large value, effectively decoupling the chains. As expected, we observe a smooth energy spectrum without gaps, and all values of the IPR become effectively zero. This result is presented in Fig. 11a.
@@
which are plotted in Fig. 11b. When these energy values are ordered from lowest to highest, they closely match the numerical spectrum in Fig. 11a. The two “tail” regions on the left and right sides of Fig. 11a contain only energies corresponding to the AA chain while the center region contains to points corresponding to both the AA and $B$ chains.
As $d$ is decreased, we treat the effects of the inter-chain coupling perturbatively Sakurai and Napolitano (2011). The diagonal matrix elements of $\hat{\tilde{H}}$ in the basis (20) are equal to zero, which means that there are no first order (or any odd order) corrections to the energy levels due to the inter-chain coupling. For even-order corrections we need to know the off-diagonal matrix elements. Using Eqs. (III) and (A) we obtain:
@@
If the incommensurability ratio $\rho$ is irrational, e.g. equal to the golden ratio $\phi$, then Eq. (24) has an infinite number of solutions for $k_{B}$ at any given $k_{A}$. To show this, let us introduce $\nu_{A}=k_{A}a/2\pi$ and $\nu_{B}=k_{B}b/2\pi$, which satisfy $0\leq\nu_{A,B}<1$. Then, we obtain from Eq. (24):
@@
In the second line here we used the fact that at $L\gg 1$ the function $F(\Theta)$ is sharply peaked near $\Theta=0$, which allows one to replace $\sin(\Theta/2)\to\Theta/2$ and extend the integration limits to infinities. Then the integral can be calculated analytically Gradshteyn and Ryzhik (2000), with $K_{1}(x)$ being the modified Bessel function of the second kind. Therefore, we obtain for the matrix elements (23):
@@
Assuming that the separation between the chains is sufficiently large to make the perturbative treatment of the inter-chain hopping quantitatively valid, one can use the large-$x$ asymptotics $K_{1}(x)\simeq\sqrt{\pi/2x}\,e^{-x}$ Abramowitz and Stegun (1970). The exponential decay of $\Phi(Q)$ means that the matrix elements rapidly decrease as $|Q|$ increases. According to Eq. (A), this constrains the possible values of the Umklapp numbers $n_{A}$ and $n_{B}$. Namely, by far the largest matrix elements correspond to $n_{A},n_{B}=0$ or 1-1, and the next largest (and considerably smaller) – to $n_{A},n_{B}=1$ or 2-2. This is illustrated in Fig. 12b, where the magnitudes $\Phi(k+Gn)$ are plotted as functions of momentum $k$ in the BZ $0\leq k<G$, for different values of $n$ (here $k=k_{A}$ or $k_{B}$, $G=2\pi/a$ or $2\pi/b$, and $n=n_{A}$ or $n_{B}$).
@@
We can now calculate the corrections to the $A$-chain dispersion. The second order correction takes the form
@@
The spectral gaps shown in Figs. 4 and 5 originate from the intra-band degeneracies. To see this, we bias one chain, say the $B$ chain, by adding a constant on‑site potential term to $\hat{H}_{B}$. For a sufficiently strong bias, the dispersion relations in the AA and $B$ chains no longer overlap and the condition (30) is never satisfied, yet the gaps persist in the energy spectrum according to our numerical results. To capture this effect, we must include the next nonzero term in the perturbation theory, namely the fourth‑order energy correction. Using the procedure for deriving the higher-order perturbative corrections developed in Refs. Huby (1961); Silverstone and Holloway (1970) we obtain:
@@
where $(\mathfrak{n}_{1},\mathfrak{n}_{2})=(1,-1)$ or (1,1)(-1,1). The solutions have the form $k^{\text{gap}}_{A}=-(\pi/b)\mathfrak{n}_{2}+(\pi/a)\ell$, where $\ell$ is an integer. Finally, using Eq. (1) we obtain for the energy gap locations in the BZ:
@@
Substituting here our values for $N_{A}$ and $N_{B}$, we find gaps located at $n=987$ and $1597$. This prediction agrees exactly with our numerical results in the case where the $B$ chain is biased such that $E^{(0)}_{B,k_{B}}\gg E^{(0)}_{A,k_{A}}$. The lower part of the spectrum corresponding to the $A$-chain dispersion modified by the coupling to the $B$ chain is displayed in Fig. 13. The upper part of the spectrum, which originates from the biased $B$ chain, is pushed up to higher energies and is not shown.

### v2

@@ V Conclusions
Previous work [44] has studied mobility edges in 1D models with explicitly imposed incommensurate modulation of the off-diagonal hopping. In that model, the localized states are exponentially localized and the localization transition is accompanied by an intermediate critical phase. In contrast, in our model the incommensurate hopping modulation arises naturally from the geometry of two coupled periodic lattices, without additional site-dependent parameters. With this different physical mechanism, we do not observe an intermediate critical phase.
@@
Experimentally, polariton setups have been used to emulate tight-binding models [6, 65], including realizations of the AA model as demonstrated in Ref. [23]. These setups utilize a planar semiconductor micro-cavity placed between two Bragg reflectors, where coupling between photons and electron–hole pairs creates exciton–polaritons. The latter represent the particles in a tight-binding model (these particles are neutral bosons). The micro-cavity is patterned into an array of identical etched micropillars, each supporting a single localized polariton mode and thus playing the role of the tight-binding lattice sites. Exponential decay of the hopping amplitudes arises naturally when the energy of the polariton mode is lower than the barrier created by the etched regions, with the decay length determined by the etching depth, which makes it possible to control hopping strength. These systems can be used to check predictions of our model, as they allow direct measurement of the wave function and can be extended to probe multiple coupled tight-binding chains with exponentially decaying hopping. Furthermore, one can synthesize an effective gauge potential in the polariton setups [41, 5, 66], making it possible to test our predictions about the effects of a magnetic field on the localization in incommensurate chains.
Our work admits several natural theoretical extensions. For instance, one can study the impact of hopping anisotropy due to $p$-wave or $d$-wave orbitals on the localization properties, as well as the differences between exponential and Gaussian hoppings. In order to better understand the transport properties of the system, one can investigate time evolution of wave packets. It would also be interesting to explore the many-body localization [2] in our system. However, this would tremendously increase the Hilbert space dimension, which is not computationally feasible with our exact diagonalization method.
@@ Acknowledgements.
Appendix A Weakly coupled chains
@@
Let us set the inter-chain separation $d$ to a sufficiently large value, effectively decoupling the chains. As expected, we observe a smooth energy spectrum without gaps, and all values of the IPR become effectively zero. This result is presented in Fig. 15(a).
@@
which are plotted in Fig. 15(b). When these energy values are ordered from lowest to highest, they closely match the numerical spectrum in Fig. 15(a). The two “tail” regions on the left and right sides of Fig. 15(a) contain only energies corresponding to the AA chain while the center region contains to points corresponding to both the AA and $B$ chains.
As $d$ is decreased, we treat the effects of the inter-chain coupling perturbatively [51]. The diagonal matrix elements of $\hat{\tilde{H}}$ in the basis (23) are equal to zero, which means that there are no first-order (or any odd order) corrections to the energy levels due to the inter-chain coupling. For even-order corrections we need to know the off-diagonal matrix elements. Using Eqs. (III) and (A) we obtain:
@@
If the incommensurability ratio $\rho$ is irrational, e.g., equal to the golden ratio $\phi$, then Eq. (27) has an infinite number of solutions for $k_{B}$ at any given $k_{A}$. To show this, let us introduce $\nu_{A}=k_{A}a/2\pi$ and $\nu_{B}=k_{B}b/2\pi$, which satisfy $0\leq\nu_{A,B}<1$. Then, we obtain from Eq. (27):
@@
In the second line here we used the fact that at $L\gg 1$ the function $F(\Theta)$ is sharply peaked near $\Theta=0$, which allows one to replace $\sin(\Theta/2)\to\Theta/2$ and extend the integration limits to infinities. Then the integral can be calculated analytically [25], with $K_{1}(x)$ being the modified Bessel function of the second kind. Therefore, we obtain for the matrix elements (26):
@@
Assuming that the separation between the chains is sufficiently large to make the perturbative treatment of the inter-chain hopping quantitatively valid, one can use the large-$x$ asymptotics $K_{1}(x)\simeq\sqrt{\pi/2x}\,e^{-x}$ [3]. The exponential decay of $\Phi(Q)$ means that the matrix elements rapidly decrease as $|Q|$ increases. According to Eq. (A), this constrains the possible values of the Umklapp numbers $n_{A}$ and $n_{B}$. Namely, by far the largest matrix elements correspond to $n_{A},n_{B}=0$ or 1-1, and the next largest (and considerably smaller) – to $n_{A},n_{B}=1$ or 2-2. This is illustrated in Fig. 16(b), where the magnitudes $\Phi(k+Gn)$ are plotted as functions of momentum $k$ in the BZ $0\leq k<G$, for different values of $n$ (here $k=k_{A}$ or $k_{B}$, $G=2\pi/a$ or $2\pi/b$, and $n=n_{A}$ or $n_{B}$).
@@
We can now calculate the corrections to the $A$-chain dispersion. The second-order correction takes the form
@@
The spectral gaps shown in Figs. 4 and 5 originate from the intra-band degeneracies. To see this, we bias one chain, say the $B$ chain, by adding a constant on‑site potential term to $\hat{H}_{B}$. For a sufficiently strong bias, the dispersion relations in the AA and $B$ chains no longer overlap and the condition (33) is never satisfied, yet the gaps persist in the energy spectrum according to our numerical results. To capture this effect, we must include the next nonzero term in the perturbation theory, namely the fourth‑order energy correction. Using the procedure for deriving the higher-order perturbative corrections developed in Refs. [29, 55] we obtain:
@@
where $(\mathfrak{n}_{1},\mathfrak{n}_{2})=(1,-1)$ or (1,1)(-1,1). The solutions have the form $k^{\text{gap}}_{A}=-(\pi/b)\mathfrak{n}_{2}+(\pi/a)\mathfrak{n}$, where $\mathfrak{n}$ is an integer. Finally, using Eq. (1) we obtain for the energy gap locations in the BZ:
@@
Figure 18: (a) Numerical spectrum of the model (4), with $N_{A}=2584$, $N_{B}=1597$, and $a=1$, as $d$ is varied. (b) Numerical spectrum of the same model with the $B$ chain biased by a uniform potential, showing only the $A$-chain band modified by coupling to the $B$ chain.
@@
Substituting here our values for $N_{A}$ and $N_{B}$, we find gaps located at $n=987$ and $1597$. This prediction agrees exactly with our numerical results in the case where the $B$ chain is biased such that $E^{(0)}_{B,k_{B}}\gg E^{(0)}_{A,k_{A}}$. The lower part of the spectrum corresponding to the $A$-chain dispersion modified by the coupling to the $B$ chain is displayed in Fig. 17. The upper part of the spectrum, which originates from the biased $B$ chain, is pushed up to higher energies and is not shown. To illustrate the evolution of spectral gaps as the inter-chain distance is varied, in Fig. 18 we plot the numerical energy spectrum as a function of $d$.